OpenStudy (anonymous):

A teacher is making a multiple choice quiz. She wants to give each student the same questions, but have each student's questions appear in a different order. If there are twenty-seven students in the class, what is the least number of questions the quiz must contain?

OpenStudy (bibby):

pretend there are only 5 questions. - - - - - there are 5 ways she can give the first one, 4 the second, etc. 5*4*3*2*1 I believe

OpenStudy (anonymous):

With one question there's only 1 permutation. 2 = 2 perms 3 = 3 * 2 * 1 = 6 4 = 4 * 3 * 2 * 1 = 24 5 = 120 permutations so would need 5 questions to give 27 students a different order.